# Federated Variational Inference for Bayesian Mixture Models

### v1

@@ I.5 Number of Batches
Appendix J Additional Simulation Results
J.1 Frequency of Merge/Delete Moves
Tables comparing wall-clock time, final log-ELBO and ARI of the resulting clustering structures in Simulations 1.1, 1.2 and 1.3 are shown below (Tables 4, 5, 6).
Table 4: Table comparing the mean wall-clock time, mean final log-ELBO, mean ARI and mean number of clusters of the final labelling structure across all runs in Simulation 1.1 with confidence intervals for time and ARI (+- 1.96 x standard deviation).
Laps Time (s) log-ELBO ARI Clusters
0 3.46 [3.43, 3.49] -25357 0.856 [0.852, 0.860] 5.01
1 6.35 [6.16, 6.55] -25354 0.858 [0.854, 0.862] 5.00
2 5.88 [5.75, 6.01] -25354 0.858 [0.854, 0.862] 5.00
5 6.75 [6.65, 6.85] -25354 0.858 [0.854, 0.862] 5.00
10 8.68 [8.57, 8.80] -25354 0.858 [0.854, 0.862] 5.00
10000 14.0 [13.2, 14.7] -25617 0.823 [0.818, 0.828] 15.3
Table 5: Table comparing the mean wall-clock time, mean final log-ELBO, mean ARI and mean number of clusters of the final labelling structure across all runs in Simulation 1.2 with confidence intervals for time and ARI (+- 1.96 x standard deviation).
Laps Time (s) log-ELBO ARI Clusters
0 10.9 [10.8, 11.0] -81256 0.959 [0.957, 0.961] 7.65
1 29.1 [26.4, 31.8] -81245 0.962 [0.960, 0.963] 7.63
2 23.4 [21.7, 25.1] -81238 0.962 [0.961, 0.963] 7.70
5 25.4 [24.1, 26.7] -81236 0.963 [0.962, 0.964] 7.84
10 31.0 [30.0, 32.0] -81260 0.962 [0.961, 0.963] 8.54
10000 47.5 [45.0, 50.0] -81721 0.940 [0.938, 0.942] 18.0
Table 6: Table comparing the mean wall-clock time, mean final log-ELBO, mean ARI and mean number of clusters of the final labelling structure across all runs in Simulation 1.3 with confidence intervals for time and ARI (+- 1.96 x standard deviation).
Laps Time (s) log-ELBO ARI Clusters
0 29.1 [28.8, 29.4] -164099 0.953 [0.952, 0.955] 10.0
1 57.1 [54.8, 59.5] -164090 0.954 [0.953, 0.955] 10.0
2 52.1 [50.7, 53.5] -164095 0.954 [0.952, 0.955] 10.0
5 61.8 [60.5, 63.1] -164100 0.954 [0.952, 0.955] 10.2
10 76.5 [74.7, 78.3] -164173 0.953 [0.951, 0.954] 12.0
10000 122 [112, 113] -164675 0.943 [0.941, 0.944] 22.7
@@
Figure S1: Graphs showing the ELBO vs wall-clock time for the algorithm for two simulated datasets across all 20 initialisations for each of the different ”laps” in Simulations 1.2 and 1.3. The mean end time across all initialisations is shown as well as an approximate 95 confidence interval.
We can clearly see from Figure S1 that a) the original variational model clearly takes far longer than models with merge/delete b) there are steep improvements in ELBO near the start with more frequent merge/delete moves. However, for example, ‘laps = 1’ is near convergence very early on in both plots, but the algorithm takes longer to actually converge and stop.
@@
Figure S2: Plot comparing final number of clusters in clustering models in simulations 1.1-1.3.
J.1.1 Variable Selection
Tables comparing wall-clock time, final log-ELBO, ARI and number of clusters of the resulting clustering structures in Simulations 1.4 and 1.5 are shown below (Tables 7, 8). We still see a speed up in terms of wall-clock time in these simulations, but see that there is more inconsistency with solutions with noisier data, especially as we increase the number of merge/delete moves, and the mean ARI for merge/delete models is generally lower when we incorporate more merge/delete moves. The number of clusters at convergence is also lower than the number we expect. We see from Figure S3 that it seems to be that merge/delete moves lead to less consistent results in the variable selection setting, but most of the best-performing models in terms of ARI are from MerDel models. As noted later (Section J.5.1), fewer clusters and lower ARIs in MerDel models could be an indication that some ‘true’ clusters could have been merged in the MerDel structures, if the ‘true’ clusters share similar characteristics.
However, the mean ELBO achieved by the resulting clustering models at convergence is still higher for all merge/delete models compared to the model with no merge/delete moves. This suggests that more care needs to be taken with the frequency of merge/delete moves used when including variable selection - it is likely that our merge/delete models are sometimes, for example, jumping too quickly to a model which has marked more variables as irrelevant and has fewer clusters, but does have a higher ELBO. MerDel models are still able to achieve some improved results, but these are more inconsistent.
Figure S3: Scatter plot comparing ARIs achieved by each model across all datasets and initialisations in Simulations 1.4 and 1.5.
Table 7: Table comparing the mean wall-clock time, mean final log-ELBO, mean ARI and mean number of clusters of the final labelling structure across all runs in Simulation 1.4 with confidence intervals for time and ARI (+- 1.96 x standard deviation).
Laps Time (s) log-ELBO ARI Clusters
0 16.9 [16.6, 17.2] -41913 0.692 [0.674, 0.710] 7.53
1 91.7 [84.0, 99.3] -41848 0.731 [0.714, 0.747] 7.86
2 75.6 [69.8, 81.4] -41838 0.753 [0.739, 0.767] 8.10
5 74.4 [70.0, 78.7] -41823 0.773 [0.759, 0.787] 8.48
10 92.7 [88.4, 96.9] -41831 0.792 [0.780, 0.804] 9.16
10000 128 [118, 138] -42372 0.782 [0.774, 0.791] 19.3
Table 8: Table comparing the mean wall-clock time, mean final log-ELBO, mean ARI and mean number of clusters of the final labelling structure across all runs in Simulation 1.5 with confidence intervals for time and ARI (+- 1.96 x standard deviation).
Laps Time (s) log-ELBO ARI Clusters
0 24.0 [23.7, 24.4] -61781 0.561 [0.540, 0.581] 6.60
1 137 [125, 149] -61664 0.625 [0.605, 0.645] 7.29
2 110 [101, 118] -61640 0.651 [0.631, 0.670] 7.67
5 112 [107, 118] -61622 0.679 [0.661, 0.696] 8.12
10 149 [143, 156] -61612 0.710 [0.695, 0.725] 8.90
10000 213 [199, 227] -62204 0.730 [0.721, 0.739] 19.3
We additionally used $F_{1}$ scores to compare the quality of variable selection. Having merge/delete moves tends to improve the quality of the variable selection as quantified by $F_{1}$ scores compared to the model with no merge/delete moves. This is illustrated in Table 9 with results from Simulation 1.4 shown. The gains are made especially in terms of identifying irrelevant, noisy variables, and laps 2 and 5 get the top $F_{1}$ scores of 0.977, indicating extremely good feature selection performance. Overall, although there is some desire for improvement in accuracy for the clustering structure when implementing variable selection, we see that merge and delete moves make improvements in identifying relevant clustering variables.
@@
J.2 Global Merge Simulations - Additional Results
Table 10: ‘Global Merge Simulations’ results. We report the median and lower/upper quantiles across all 10 ‘shuffles’/initialisations and all 10 synthetic datasets. FedMerDel/G is the greedy search, FedMerDel/R is the random search for FedMerDel. Note separate independent datasets were generated for $N$ equal but different numbers of batches/cores, accounting for slight differences in ARI.
@@
N Model Batches/Cores ARI Clusters Time (s)
20000 full 5 0.943 [0.933, 0.950] 12 [11.8, 12] 106 [92.8, 130]
par 5 0.943 [0.930, 0.950] 12 [11, 12] 37.0 [33.4, 40.6]
FedMerDel/G 5 0.920 [0.912, 0.933] 12 [12, 12] 33.1 [28.9, 37.2]
FedMerDel/R 5 0.920 [0.912, 0.932] 12 [12, 12] 33.2 [29.1, 37.4]
50000 full 5 0.947 [0.941, 0.954] 12 [12, 12] 467 [414, 554]
par 5 0.948 [0.942, 0.954] 12 [12, 12] 176 [163, 196]
FedMerDel/G 5 0.945 [0.939, 0.949] 13 [13, 14] 114 [106, 123]
FedMerDel/R 5 0.945 [0.939, 0.948] 12 [12, 12] 114 [106, 123]
50000 full 10 0.957 [0.955, 0.958] 12 [12, 12] 465 [432, 512]
par 10 0.956 [0.955, 0.958] 12 [12, 12] 126 [121, 142]
FedMerDel/G 10 0.951 [0.946, 0.954] 13 [13, 14] 70.6 [66.2, 75.8]
FedMerDel/R 10 0.951 [0.946, 0.954] 12 [12, 13] 71.3 [66.8, 76.6]
100000 full 5 0.955 [0.944, 0.961] 12 [12, 12] 956 [861, 1116]
par 5 0.954 [0.944, 0.961] 12 [12, 12] 338 [310, 380]
FedMerDel/G 5 0.954 [0.943, 0.960] 13 [12, 13] 260 [231, 287]
FedMerDel/R 5 0.954 [0.943, 0.960] 12 [12, 12] 260 [232, 287]
100000 par 10 0.949 [0.941, 0.959] 12 [12, 12] 249 [231, 284]
FedMerDel/G 10 0.948 [0.938, 0.952] 13 [13, 14] 131 [123, 143]
FedMerDel/R 10 0.948 [0.938, 0.952] 12 [12, 13] 132 [124, 144]
200000 par 10 0.952 [0.947, 0.958] 12 [12, 12] 482 [446, 536]
FedMerDel/G 10 0.951 [0.946, 0.957] 13 [13, 14] 269 [242, 303]
FedMerDel/R 10 0.951 [0.946, 0.957] 12.5 [12, 13] 270 [243, 305]
500000 par 10 0.950 [0.943, 0.951] 12 [12, 12] 1121 [1056, 1303]
FedMerDel/G 10 0.949 [0.942, 0.951] 13 [13, 14] 685 [642, 793]
FedMerDel/R 10 0.949 [0.942, 0.951] 12.5 [12, 13] 688 [646, 797]
1000000 par 20 0.948 [0.943, 0.953] 12 [12, 12] 2119 [2031, 2204]
FedMerDel/G 20 0.948 [0.943, 0.953] 13.5 [13, 14] 899 [825, 984]
FedMerDel/R 20 0.948 [0.943, 0.953] 14 [13, 14] 912 [840, 1007]
J.2.1 K=6 simulations
Similar results were seen as with the simulation scenarios with $K=12$, where the gains in time efficiency from using FedMerDel instead of parallelised MerDel increased as $N$ increased.
@@
Table 11: Results from ‘Global Merge Simulations’ with $K=6$ clusters, where all datasets have $P=100$ covariates, and are split into 20 batches or parallelised over 20 cores. We initialised with 10 clusters in all MerDel cases. We compare FedMerDel to parallelised MerDel. We report the median and lower/upper quantiles across 5 ‘shuffles’/initialisations for 5 synthetic datasets. Results are for the random search; greedy searches differed insignificantly.
N Model ARI Clusters Time (s)
100000 par 0.974 [0.967, 0.978] 6 [6, 6] 111 [103, 121]
FedMerDel 0.965 [0.972, 0.977] 6 [6, 6] 70.5 [66.2, 77.6]
500000 par 0.970 [0.969, 0.975] 6 [6, 6] 645 [619, 734]
FedMerDel 0.970 [0.969, 0.975] 6 [6, 6] 455 [404, 516]
J.2.2 Varying P
@@
P Model ARI Clusters Time (s)
60 par 0.782 [0.753, 0.796] 10 [10, 10] 206 [194, 231]
FedMerDel 0.780 [0.751, 0.794] 10 [10, 10] 153 [130, 169]
80 par 0.874 [0.872, 0.880] 10 [10, 10] 285 [256, 333]
FedMerDel 0.872 [0.870, 0.879] 10 [10, 10] 190 [182, 231]
120 par 0.972 [0.972, 0.973] 10 [10, 10] 316 [292, 345]
FedMerDel 0.972 [0.971, 0.972] 10 [10, 10] 208 [203, 229]
200 par 0.999 [0.999, 0.999] 10 [10, 10] 465 [453, 469]
FedMerDel 0.999 [0.999, 0.999] 10 [10, 10] 295 [291, 302]
Figure S4: Plot comparing the mean time taken by parallelised MerDel (‘par’) and FedMerDel (‘shard’) as we vary $P$, with a 95% confidence interval (mean $\pm$ 1.96 ×$\times\frac{s.d}{\sqrt{n}}$). Results for FedMerDel are with the greedy search, although there is no visible difference for the random search.
J.2.3 Global Merge with Variable Selection
Results showed similar performance in terms of ARI between FedMerDel and MerDel on the full dataset (Table 13). In these scenarios (with relatively low $N$) FedMerDel did not outperform parallelised MerDel in terms of wall-clock time, but notably was almost always able to correctly identify relevant and irrelevant variables, outperforming both other methods, which usually wrongly identified some irrelevant variables as relevant. Global merge times were less than one second in all cases.

### v2

@@ G.5 Number of Batches
Appendix H Additional Simulation Results
H.1 Frequency of Merge/Delete Moves
Incorporating merge/delete enables faster convergence in terms of wall-clock time: Generally, MerDel models are faster to converge. Figure 6 in Appendix H.1 shows that early accepted merge/delete moves allow for substantial early increases in the ELBO. However, as described in Section 4, there is a trade-off that must be made - models with laps = 1, for example, get close to convergence very fast but then take more time to fully converge. The inclusion of frequent merge/delete moves at the start of the algorithm can also lead to the algorithm jumping to a worse local optimum, reflected in slightly lower ARIs in some cases.
Better ARIs are achieved with merge/delete: Figure LABEL:merdelARIplot and Appendix H.1 show that laps = 2, 5, 10 generally yield higher ARIs, particularly in Simulations 1.1-1.3.
[…]
merdelARIplot Refer to caption
Merge/delete more accurately estimates the number of clusters: The model with no merge/delete moves tends to vastly overestimate the number of true clusters in the data due to the existence of small unnecessary clusters. MerDel mitigates this, with Figure 5 in Appendix H.1 showing that the model often is able to find the exact true number.
Table LABEL:tablesim_combined shows wall-clock time, final log-ELBO and ARI of the resulting clustering structures in Simulations 1.1, 1.2 and 1.3.
[…]
tablesim_combined Simulation Laps Time (s) log-ELBO ARI Clusters 1.1 0 3.46 [3.43, 3.49] -25357 0.856 [0.852, 0.860] 5.01 1 6.35 [6.16, 6.55] -25354 0.858 [0.854, 0.862] 5.00 2 5.88 [5.75, 6.01] -25354 0.858 [0.854, 0.862] 5.00 5 6.75 [6.65, 6.85] -25354 0.858 [0.854, 0.862] 5.00 10 8.68 [8.57, 8.80] -25354 0.858 [0.854, 0.862] 5.00 10000 14.0 [13.2, 14.7] -25617 0.823 [0.818, 0.828] 15.3 1.2 0 10.9 [10.8, 11.0] -81256 0.959 [0.957, 0.961] 7.65 1 29.1 [26.4, 31.8] -81245 0.962 [0.960, 0.963] 7.63 2 23.4 [21.7, 25.1] -81238 0.962 [0.961, 0.963] 7.70 5 25.4 [24.1, 26.7] -81236 0.963 [0.962, 0.964] 7.84 10 31.0 [30.0, 32.0] -81260 0.962 [0.961, 0.963] 8.54 10000 47.5 [45.0, 50.0] -81721 0.940 [0.938, 0.942] 18.0 1.3 0 29.1 [28.8, 29.4] -164099 0.953 [0.952, 0.955] 10.0 1 57.1 [54.8, 59.5] -164090 0.954 [0.953, 0.955] 10.0 2 52.1 [50.7, 53.5] -164095 0.954 [0.952, 0.955] 10.0 5 61.8 [60.5, 63.1] -164100 0.954 [0.952, 0.955] 10.2 10 76.5 [74.7, 78.3] -164173 0.953 [0.951, 0.954] 12.0 10000 122 [112, 113] -164675 0.943 [0.941, 0.944] 22.7
Table 5: Performance comparison across Simulations 1.1, 1.2, and 1.3: mean wall-clock time, log-ELBO, ARI and number of clusters with confidence intervals for time and ARI (±1.96 × standard deviation).
Figure 5: Plot comparing final number of clusters in clustering models in simulations 1.1-1.3.
@@ H.1 Frequency of Merge/Delete Moves
Figure 6: Graphs showing the ELBO vs wall-clock time for the algorithm for two simulated datasets across all 20 initialisations for each of the different ”laps” in Simulations 1.2 and 1.3. The mean end time across all initialisations is shown, as well as an approximate 95% confidence interval.
We can clearly see from Figure 6 that a) the original variational model clearly takes far longer than models with merge/delete b) there are steep improvements in ELBO near the start with more frequent merge/delete moves. However, for example, ‘laps = 1’ is near convergence very early on in both plots, but the algorithm takes longer to actually converge and stop.
@@
H.1.1 Variable Selection
Table LABEL:tablesim45_combined shows wall-clock time, final log-ELBO and ARI of the resulting clustering structures in Simulations 1.4 and 1.5. We still see a speed up in terms of wall-clock time in these simulations, but see that there is more inconsistency with solutions with noisier data, especially as we increase the number of merge/delete moves, and the mean ARI for merge/delete models is generally lower when we incorporate more merge/delete moves. The number of clusters at convergence is also lower than expected. We see from Figure 7 that it seems to be that merge/delete moves lead to less consistent results in the variable selection setting, but most of the best-performing models in terms of ARI are from MerDel models.
Figure 7: Scatter plot comparing ARIs across all datasets and initialisations in Simulations 1.4 and 1.5.
[…]
tablesim45_combined Simulation Laps Time (s) log-ELBO ARI Clusters 1.4 0 16.9 [16.6, 17.2] -41913 0.692 [0.674, 0.710] 7.53 1 91.7 [84.0, 99.3] -41848 0.731 [0.714, 0.747] 7.86 2 75.6 [69.8, 81.4] -41838 0.753 [0.739, 0.767] 8.10 5 74.4 [70.0, 78.7] -41823 0.773 [0.759, 0.787] 8.48 10 92.7 [88.4, 96.9] -41831 0.792 [0.780, 0.804] 9.16 10000 128 [118, 138] -42372 0.782 [0.774, 0.791] 19.3 1.5 0 24.0 [23.7, 24.4] -61781 0.561 [0.540, 0.581] 6.60 1 137 [125, 149] -61664 0.625 [0.605, 0.645] 7.29 2 110 [101, 118] -61640 0.651 [0.631, 0.670] 7.67 5 112 [107, 118] -61622 0.679 [0.661, 0.696] 8.12 10 149 [143, 156] -61612 0.710 [0.695, 0.725] 8.90 10000 213 [199, 227] -62204 0.730 [0.721, 0.739] 19.3
Table 6: Performance comparison across Simulations 1.4 and 1.5: mean wall-clock time, log-ELBO, ARI and number of clusters with confidence intervals for time and ARI (±1.96 × standard deviation).
We additionally used $F_{1}$ scores to compare the quality of variable selection. Having merge/delete moves tends to improve the quality of the variable selection as quantified by $F_{1}$ scores. This is illustrated in Table 7. The gains are made especially in terms of identifying irrelevant, noisy variables, and laps 2 and 5 get the top $F_{1}$ scores of 0.977, indicating extremely good feature selection performance.
@@ H.1.1 Variable Selection
H.2 Global Merge Simulations - Additional Results
[…]
@@ H.2 Global Merge Simulations - Additional Results
globalmergeresultsfull $\mathbf{N}$ /Sim Model Batches ARI Clusters Time (s) 20000 full 5 0.943 [0.933, 0.950] 12 [11.8, 12] 106 [92.8, 130] par 5 0.943 [0.930, 0.950] 12 [11, 12] 37.0 [33.4, 40.6] FedMerDel/G 5 0.920 [0.912, 0.933] 12 [12, 12] 33.1 [28.9, 37.2] FedMerDel/R 5 0.920 [0.912, 0.932] 12 [12, 12] 33.2 [29.1, 37.4] 50000 full 5 0.947 [0.941, 0.954] 12 [12, 12] 467 [414, 554] par 5 0.948 [0.942, 0.954] 12 [12, 12] 176 [163, 196] FedMerDel/G 5 0.945 [0.939, 0.949] 13 [13, 14] 114 [106, 123] FedMerDel/R 5 0.945 [0.939, 0.948] 12 [12, 12] 114 [106, 123] 50000 full 10 0.957 [0.955, 0.958] 12 [12, 12] 465 [432, 512] par 10 0.956 [0.955, 0.958] 12 [12, 12] 126 [121, 142] FedMerDel/G 10 0.951 [0.946, 0.954] 13 [13, 14] 70.6 [66.2, 75.8] FedMerDel/R 10 0.951 [0.946, 0.954] 12 [12, 13] 71.3 [66.8, 76.6] 100000 full 5 0.955 [0.944, 0.961] 12 [12, 12] 956 [861, 1116] par 5 0.954 [0.944, 0.961] 12 [12, 12] 338 [310, 380] FedMerDel/G 5 0.954 [0.943, 0.960] 13 [12, 13] 260 [231, 287] FedMerDel/R 5 0.954 [0.943, 0.960] 12 [12, 12] 260 [232, 287] 100000 par 10 0.949 [0.941, 0.959] 12 [12, 12] 249 [231, 284] FedMerDel/G 10 0.948 [0.938, 0.952] 13 [13, 14] 131 [123, 143] FedMerDel/R 10 0.948 [0.938, 0.952] 12 [12, 13] 132 [124, 144] 200000 par 10 0.952 [0.947, 0.958] 12 [12, 12] 482 [446, 536] FedMerDel/G 10 0.951 [0.946, 0.957] 13 [13, 14] 269 [242, 303] FedMerDel/R 10 0.951 [0.946, 0.957] 12.5 [12, 13] 270 [243, 305] 500000 par 10 0.950 [0.943, 0.951] 12 [12, 12] 1121 [1056, 1303] FedMerDel/G 10 0.949 [0.942, 0.951] 13 [13, 14] 685 [642, 793] FedMerDel/R 10 0.949 [0.942, 0.951] 12.5 [12, 13] 688 [646, 797] 1000000 par 20 0.948 [0.943, 0.953] 12 [12, 12] 2119 [2031, 2204] FedMerDel/G 20 0.948 [0.943, 0.953] 13.5 [13, 14] 899 [825, 984] FedMerDel/R 20 0.948 [0.943, 0.953] 14 [13, 14] 912 [840, 1007] (a) par 10 0.945 [0.944, 0.951] 12 [12, 12] 119 [110, 131] FedMerDel 10 0.942 [0.936, 0.946] 13 [12, 13] 64.5 [61.5, 70.1] (b) par 5 0.955 [0.949, 0.956] 10 [10, 10] 169 [161, 192] FedMerDel 5 0.993 [0.992, 0.994] 10 [10, 10] 131 [127, 165] (c) par 5 0.950 [0.944, 0.950] 12 [12, 12.8] 69.5 [65.8, 77.3] FedMerDel 5 0.988 [0.985, 0.990] 13 [12, 13] 53.2 [48.3, 58.3]
@@
Table 8: ‘Global Merge Simulations’ results. We report the median and lower/upper quantiles across all 10 ‘shuffles’/initialisations and all 10 synthetic datasets. FedMerDel/G is the greedy search, FedMerDel/R is the random search for FedMerDel. Note separate independent datasets were generated for NN equal but different numbers of batches/cores, accounting for slight differences in ARI. Heterogeneous simulations (a), (b), (c) are described in Appendix G.4.
H.2.1 Comparisons to Other Unsupervised Centralised Learning Methods
While k-modes, hierarchical clustering and DBSCAN were fast and clearly superior in terms of computational efficiency (eg. k-modes took around 25 seconds on datasets of size $N=50,000$, all performed very poorly with ARIs close to 0 compared to the true simulated labels, suggesting these were barely better than random cluster assignment. These methods also face limitations; they often require manual specification of the number of clusters, have limited interpretability and struggle with high dimensionality.
Using the poLCA R package on a simulated dataset with N = 50,000, LCA achieved a mean ARI of 0.867, underperforming both FedMerDel and parallelised MerDel. Furthermore, this required around 10 minutes per model. LCA requires users to fit multiple models and compare with criteria such as BIC to determine the number of clusters. Crucially, this method is not directly applicable to FL scenarios.
H.2.2 Varying P
@@
[…]
globalmergeresultsP $\mathbf{N}$ Model ARI Clusters Time (s) 60 par 0.782 [0.753, 0.796] 10 [10, 10] 206 [194, 231] FedMerDel 0.780 [0.751, 0.794] 10 [10, 10] 153 [130, 169] 80 par 0.874 [0.872, 0.880] 10 [10, 10] 285 [256, 333] FedMerDel 0.872 [0.870, 0.879] 10 [10, 10] 190 [182, 231] 120 par 0.972 [0.972, 0.973] 10 [10, 10] 316 [292, 345] FedMerDel 0.972 [0.971, 0.972] 10 [10, 10] 208 [203, 229] 200 par 0.999 [0.999, 0.999] 10 [10, 10] 465 [453, 469] FedMerDel 0.999 [0.999, 0.999] 10 [10, 10] 295 [291, 302]
@@
Figure 8: Plot comparing the mean time taken by parallelised MerDel (‘par’) and FedMerDel (‘shard’) as we vary $P$, with a 95% confidence interval (mean $\pm$ 1.96 $\times\frac{s.d}{\sqrt{n}}$). Results for FedMerDel are with the greedy search.
H.2.3 Less Separable Clusters
In both simulation scenarios with the alternative generating distribution to allow for less cluster separability, results still remained robust as before, where ARIs remained close to 0.99. Furthermore, in the second scenario where 75% of covariates were similar across clusters 11 and 12, in all cases, FedMerDel separated the observations from these clusters into distinct clusters.
H.2.4 Global Merge with Variable Selection
Results showed similar performance in terms of ARI between FedMerDel and MerDel on the full dataset (Table LABEL:globalmergeresultsvs). In these scenarios (with relatively low $N$) FedMerDel notably was almost always able to correctly identify relevant and irrelevant variables, outperforming other methods, which usually wrongly identified some irrelevant variables as relevant. Global merges took less than one second in all cases.
[…]
globalmergeresultsvs $\mathbf{N}$ Model ARI Clusters Time (s) Rel Irrel 20000 full 0.865 [0.862, 0.868] 15 [15, 15] 782 [687, 835] 80 [80, 80] 18 [17, 18] par 0.865 [0.862, 0.868] 15 [15, 15] 207 [187, 228] 80 [80, 80] 17 [16, 17] FedMerDel 0.850 [0.844, 0.857] 15 [15, 15] 223 [212, 242] 80 [80, 80] 20 [20, 20] 50000 full 0.885 [0.876, 0.885] 10 [10, 10] 1607 [1493, 1767] 75 [75, 75] 23 [22, 24] par 0.885 [0.876, 0.886] 10 [10, 10] 462 [440, 527] 75 [75, 75] 23 [22, 24] FedMerDel 0.881 [0.873, 0.883] 10 [10, 10] 630 [563, 672] 75 [75, 75] 25 [25, 25]